OpenStudy (anonymous):

Determine the 5th term of an infinite geometric series if S(infinite)=18 and r=2/3

OpenStudy (mertsj):

\[Sum = \frac{a}{1-r}\] \[18=\frac{a}{1-\frac{2}{3}}\] \[54=a\] \[a _{n}=a _{1}r ^{n-1}\] \[a _{5}=54(\frac{2}{3})^{4}=54(\frac{16}{81})=\frac{32}{3}\]